A stockbroker has money in three accounts. The interest rates on the three accounts are 6%, 8%, and 10%. If she has twice as much money invested at 8% as she has invested at 6%, three times as much at 10% as she has at 6%, and the total interest for the year is $208, how much is invested at each rate?
Answer by mananth(16946)

6% ---------x
8%---------2x
10%---------3x
6%*x+8%*2x+10%*3x=208
multiply equation by 100
6x+16x+30x=20800
52x= 20800
x= 20800/52
x=400
you can calculate the others by plugging